OpenStudy (anonymous):

How do I write the equation of the circle in standard form, and determine the center and radius if it exists? for: 2x^2 + 2y^2 + 8x - 12y - 8 = 0

OpenStudy (amistre64):

complete the squares if need be

OpenStudy (amistre64):

2(x^2 +4x) + 2(y^2 -6y) = 8 2(x^2 +4x +4)-8 + 2(y^2 -6y+9)-18 = 8

OpenStudy (amistre64):

2(x+2)^2 + 2(y-3)^2 = 8+8+18 = 34

OpenStudy (amistre64):

the center is the opposites of the x and y modifiers... C=(-2,3)

OpenStudy (amistre64):

r^2 = 34 r = sqrt(34)

OpenStudy (amistre64):

or do we divide it all by 2? first?

OpenStudy (amistre64):

r=sqrt(17) prolly

OpenStudy (anonymous):

Not easy, so do they both get divided by 2 then?

OpenStudy (amistre64):

everything gets divided by 2...everything

OpenStudy (anonymous):

Okay,

OpenStudy (amistre64):

on a side note, we could divide everying by 34 to get: (x+2)^2 (y-3)^2 ------- + ------- = 1 ; which is the same equation 17 17

OpenStudy (amistre64):

so I would stick with r = sqrt(17)

OpenStudy (anonymous):

Okay, thank you

OpenStudy (amistre64):

youre welcome :)